# Optimal control under unknown intensity with Bayesian learning

### v1

@@ 3.1 Space regularity
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}J_{1}^{\gamma}(t,y,z,n)+J^{\gamma}_{2}(t,y,z,n),$
@@
where under $\mathbb{Q}$, $\Lambda$ is a random variable with law $\mathbb{P}(\Lambda\in\mathrm{d}\lambda)=\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda)/\Phi_{\mu}(n,z)$ and $(N_{t})$ is a standard Poisson process with rate $1$ independent of $\Lambda$.
@@
    The prior distribution is compactly supported: $\exists\lambda_{\max}>0$:
@@
    The function $g\in C^{1}(\mathbb{R};\mathbb{R}_{+})$ with $\lVert g\rVert_{\infty}+\lVert\frac{g^{\prime}}{g}\rVert_{\infty}<\infty$.
@@
    The function $b\in C^{1}(\mathbb{R})$ with $\lVert b^{\prime}\rVert_{\infty}<\infty$.
The main result of this section is the following.
@@
Under Assumption 3.1, it holds that there exists a constant $C_{T}$ such that for all $n\geq 0$ and for all $x,x^{\prime}\in\mathbb{R}$ and all $z,z^{\prime}\geq 0$:
@@
We now give the proof of this result. To simplify the notations, we assume that $t=0$ and write:
@@ 3.1.1 Regularity with respect to $y$
There exists a constant $C_{T}$ such that for all $\gamma\in{\mathcal{A}}$, it holds that:
@@
We prove the result with $C_{T}:=\left(\lVert\frac{g^{\prime}}{g}\rVert_{\infty}+\lambda_{\max}T\lVert g^{\prime}\rVert_{\infty}\right)e^{T\lVert b^{\prime}\rVert_{\infty}}$. Observe that just after a jump, $Y$ is reset to zero and so the initial condition $y$ is forgotten. We write $\varphi_{s}(y)$ for the solution of the ODE $\frac{d}{ds}\varphi_{s}(y)=b(\varphi_{s}(y))+\gamma_{u}$ with initial condition $\varphi_{0}(y)=y$. Assume first that $N_{T}=0$. Then
@@
As $\partial_{y}\varphi_{s}(y)=\exp\left(\int_{0}^{s}b^{\prime}(\varphi_{u}(y))\mathrm{d}u\right)\leq\exp(T\lVert b^{\prime}\rVert_{\infty})$, the result holds.
Assume now that $N_{T}>0$. Then, provided that $\tau_{1}$ is the time of the first jump of $N$, we have:
@@
We deduce the result using our assumptions on $g$ and $b$. ∎
@@
There exists a constant $C_{T}$ such that for all $\gamma\in{\mathcal{A}}$,
@@
Assume first that $y^{\prime}>y$. Then the result follows by Grönwall’s lemma. If now $y^{\prime}<y$, we set $\phi(s)=L^{y-s,\gamma}_{T}(\lambda)$. We have $\phi^{\prime}(s)=-\partial_{y}L^{y-s,\gamma}_{T}(\lambda)\leq C_{T}\phi(s)$. Therefore, by Grönwall’s lemma, $\phi(s)\leq e^{C_{T}s}\phi(0).$ Choosing $s=y-y^{\prime}$ ends the proof. ∎
@@
There is another constant $C_{T}$ such that for any $\gamma\in{\mathcal{A}}$, for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$:
@@
Without loss of generality, assume that $y+1>y^{\prime}>y$. Then we have $L^{y^{\prime},\gamma}_{T}(\lambda)-L^{y,\gamma}_{T}(\lambda)\leq(e^{C_{T}(y^{\prime}-y)}-1)L^{y,\gamma}_{T}(\lambda)\leq C_{T}e^{C_{T}}|y-y^{\prime}|L^{y,\gamma}_{T}(\lambda)$. Similarly, $L^{y,\gamma}_{T}(\lambda)-L^{y^{\prime},\gamma}_{T}(\lambda)\leq C_{T}e^{C_{T}}|y-y^{\prime}|L^{y^{\prime},\gamma}_{T}(\lambda)$. Altogether, the result holds. ∎
@@
As $L^{y^{\prime},\gamma}_{T}(\lambda)\leq e^{C_{T}}L^{y,\gamma}_{T}(\lambda)$, we deduce that there exists another constant $C_{T}$ such that for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$,
@@
There exists a constant $C_{T}$ such that for any $\gamma\in{\mathcal{A}}$, it holds that for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$:
@@
We have $A\leq C_{T}|y-y^{\prime}|\mathbb{E}_{\mathbb{Q}}L^{\gamma,y}_{t,T}(\Lambda)\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})=C_{T}|y-y^{\prime}|J^{\gamma}_{2}(t,y,z,n)$ and:
@@
We used that $\sup_{n\in\mathbb{N}}\lVert\partial_{z}\Psi_{\mu}(n,\cdot)\rVert<\infty$. We deduce the result. ∎
Altogether, we deduce that
@@
Let $y,y^{\prime}\in\mathbb{R}$ be fixed. Consider $\epsilon>0$ and $\gamma$ an $\epsilon$-optimal control in the sense that
@@
Sending $\epsilon\downarrow 0$, we deduce that
@@
In addition, note that
@@
Exchanging the role of $y$ and $y^{\prime}$ we deduce that
@@ 3.1.2 Regularity with respect to $z$
The regularity with respect to $z$ is proven by similar arguments. Recall that
@@
There exists constants $C,\eta>0$ such that for all $n\in\mathbb{N}$:
@@
    $\forall z,z^{\prime}\geq 0$, $\lvert\Phi_{\mu}(n,z)-\Phi_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|\Phi_{\mu}(n,z)$.
@@
    $\forall z,z^{\prime}\geq 0$, $\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|[m_{\mu}(n,z)+m_{\mu}(n,z^{\prime})]$.
@@
    $\forall z,z^{\prime}\geq 0$ with $|z-z^{\prime}|<\eta$, $\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|m_{\mu}(n,z)$.
@@
We start with the first point. We have
@@
As $Supp(\mu)\subset[0,\lambda_{\max}]$, the result holds. For the second point, we have
@@
Using the first point, we deduce that the second inequality holds. Finally, we verify that provided $|z-z^{\prime}|\leq\eta$, $\eta$ small enough, it holds that
@@
As $\lvert\Phi_{\mu}(n,z^{\prime})-\Phi_{\mu}(n,z)\rvert\leq\frac{1}{2}\Phi_{\mu}(n,z)$ provided that $\lvert z-z^{\prime}\rvert$ is small enough, we deduce the last statement. ∎
As on $\mathbb{Q}$, the Poisson process $N$ is independent of $\Lambda$, we have
@@
We deduce from the last point of the previous Lemma that for all $z,z^{\prime}\geq 0$, $\lvert z-z^{\prime}\rvert\leq\eta$,
@@
We deduce as previously that
@@
This ends the proof of Proposition 3.2.
@@ 3.2 A first dynamic programming principle
Recall that the value function is defined by
$v(t,y,z,n)$ $:=\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\left[\int_{t}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+\kappa\Psi_{\mu}(n+N_{T}-N_{t},Z_{T}^{t,y,z,\gamma})\right]$
$=\inf_{\gamma\in{\mathcal{A}}}J^{\gamma}(t,y,z,n).$
@@
It holds that for all $s\in[t,T]$,
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}$ $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left[\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right].$
We now detail the proof of this result. Recall that for $\gamma\in{\mathcal{A}}$,
@@
We start with the following key lemma.
@@
Let $S$ be a $({\mathcal{F}}^{N}_{t})$-stopping time such that $a.s.$, $S\in[t,T]$. It holds that $a.s$,
$\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\int_{S}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u\mid{\mathcal{F}}^{N}_{S}\right]$
$\quad=\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle J^{\gamma}_{1}(S,Y^{t,y,\gamma}_{S},Z^{t,y,z,\gamma}_{S},n+N_{S}-N_{t}),$
and
$\mathbb{E}_{\mathbb{Q}}[\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})~{}|~{}{\mathcal{F}}^{N}_{S}]$
$\quad=\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle J^{\gamma}_{2}(S,Y^{t,y,\gamma}_{S},Z^{t,y,z,\gamma}_{S},n+N_{S}-N_{t}).$
@@
First, we note that for all $\lambda\geq 0$:
$a.s,\quad L^{y,\gamma}_{t,T}(\lambda)=L^{y,\gamma}_{t,S}(\lambda)L^{\gamma,Y^{t,y,\gamma}_{S}}_{S,T}(\lambda).$
We deduce from (2.7) that
$\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\langle L^{Y^{t,y,\gamma}_{S},\gamma}_{S,T},m_{\mu}(n+N_{S}-N_{t},Z^{t,y,z,\gamma}_{S}\rangle=\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle.$
We now give the proof of the second equality. Take $A\in{\mathcal{F}}^{N}_{S}$. To simplify the notations, write:
$y_{S}:=Y^{t,y,\gamma}_{S},\quad z_{S}:=Z^{t,y,z,\gamma}_{S},\quad n_{S}:=n+N_{S}-N_{t}.$
We also set:
$\Delta Z_{S,T}:=Z_{T}^{S,y_{S},z_{S},\gamma}-z_{S}=\int_{S}^{T}g(Y^{t,y,\gamma}_{u})\mathrm{d}u.$
We have
@@
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\langle L^{y_{S},\gamma}_{S,T},m_{\mu}(n_{S},z_{S})\rangle\mathbbm{1}_{A}\Psi_{\mu}(n_{S}+N_{T}-N_{S},z_{S}+\Delta Z_{S,T})$
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\mathbbm{1}_{A}$
$\qquad~{}~{}~{}~{}~{}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y_{S},\gamma}_{S,T},m_{\mu}(n_{S},z_{S})\rangle\Psi_{\mu}(n_{S}+N_{T}-N_{S},z_{S}+\Delta Z_{S,T})~{}|~{}{\mathcal{F}}^{N}_{S}\right]$
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\mathbbm{1}_{A}J^{\gamma}_{2}(S,y_{S},z_{S},n_{S}).$
The obtain the last identity, we used the strong Markov property satisfied by $(Y^{t,y,\gamma}_{s},n+N_{s}-N_{t},Z^{t,y,z,\gamma}_{s})$. As the equality is valid for any $A\in{\mathcal{F}}^{N}_{S}$, the result is proved. ∎
We now detail the proof of Proposition 3.9, which follow from standard arguments. Recall that $J^{\gamma}(t,y,z,n)=J^{\gamma}_{1}(t,y,z,n)+\kappa J^{\gamma}_{2}(t,y,z,n)$. Using Lemma 3.10 with a deterministic $S=s\in[t,T]$, we have
@@
We deduce that
$v(t,y,z,n)\geq\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{x,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left\{\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right\}\right].$
For the other inequality, fix $\epsilon>0$ and $s\in(0,T)$. Using the regularity of $J^{\gamma}$ and $v$ proven in Section 3.1, there exists $\alpha>0$ such that for all $n\in\mathbb{N}$, there is a partition $(B_{i}^{n})_{i\in\mathbb{N}}$ of $\mathbb{R}\times\mathbb{R}_{+}$ with centers $(y_{i}^{n},z_{i}^{n})\in B_{i}^{n}$ such that: for all $(n,i)\in\mathbb{N}^{2}$, for all $(y,z)\in B_{i}^{n}$,
@@
For each $(y_{i}^{n},z_{i}^{n},n)$, there exists an $\epsilon$-optimal control $\alpha(y_{i}^{n},z_{i}^{n},n)\in{\mathcal{A}}$:
@@
Finally, we set for all $\gamma\in{\mathcal{A}}$:
@@
where $\alpha^{i,n}:=\alpha(y_{i}^{n},z_{i}^{n},n)$ for $i,n\geq 0$ and $n_{s}:=n+N_{s}-N_{t}$. To simplify the notations, we also write $y_{s}=Y^{t,y,\gamma}_{s}$ and $z_{s}=Z^{t,y,z,\gamma}_{s}$. We have:
@@
We send $\epsilon$ to zero. Using moreover that the control $\gamma$ is arbitrary between $t$ and $s$, we deduce that

### v2

@@ 3 Dynamic programming principle and regularity
We now establish that the value function is regular and satisfies a dynamic programming principle. Recall that we work in the canonical space defined in Section 2.4. For simplicity, given $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$ and a measurable function $h:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}_{+}$, we write:
$\mathbb{E}_{\mathbb{Q}}[h(\Lambda,N)]:=\int_{\mathbb{R}_{+}}\mathbb{E}_{\mathbb{Q}}[h(\lambda,N)]\mu(\mathrm{d}\lambda).$
@@ 3.1 Space regularity
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}_{t}}J_{1}^{\gamma}(t,y,z,n)+J^{\gamma}_{2}(t,y,z,n),$
@@
and $\Lambda$ is a random variable with distribution $\mathbb{P}(\Lambda\in\mathrm{d}\lambda)=\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda)/\Phi_{\mu}(n,z)$.
@@
    The prior distribution is compactly supported: there exists $\lambda_{\max}>0$ such that:
@@
    The function $g\in C^{1}(\mathbb{R};\mathbb{R}_{+})$ satisfies $\lVert g\rVert_{\infty}+\lVert\frac{g^{\prime}}{g}\rVert_{\infty}<\infty$.
@@
    The function $b\in C^{1}(\mathbb{R})$ satisfies $\lVert b^{\prime}\rVert_{\infty}<\infty$.
The main result of this section is the following:
@@
Under Assumption 3.1, there exists a constant $C_{T}$ such that, for all $n\geq 0$, $y,y^{\prime}\in\mathbb{R}$, and $z,z^{\prime}\geq 0$:
@@
We now provide the proof of this result. To simplify notation, we assume $t=0$ and write:
@@ 3.1.1 Regularity with respect to $y$
There exists a constant $C_{T}$ such that, for all $\gamma\in{\mathcal{A}}$:
@@
[…]
Case 1: $N_{T}=0$. If there are no jumps, then:
@@
Since $\partial_{y}\varphi_{s}(y)=\exp\left(\int_{0}^{s}b^{\prime}(\varphi_{u}(y))\mathrm{d}u\right)\leq\exp(T\lVert b^{\prime}\rVert_{\infty})$, the result holds.
Case 2: $N_{T}>0$. Let $\tau_{1}$ denote the time of the first jump of $N$. Then:
@@
The result follows from the assumptions on $g$ and $b$. ∎
@@
There exists a constant $C_{T}$ such that, for all $\gamma\in{\mathcal{A}}$:
@@
Assume first that $y^{\prime}>y$. The result follows by applying Grönwall’s lemma. If now $y^{\prime}<y$, define $\phi(s)=L^{y-s,\gamma}_{T}(\lambda)$. Then:
$\phi^{\prime}(s)=-\partial_{y}L^{y-s,\gamma}_{T}(\lambda)\leq C_{T}\phi(s).$
By Grönwall’s lemma, $\phi(s)\leq e^{C_{T}s}\phi(0)$. Choosing $s=y-y^{\prime}$ completes the proof. ∎
@@
There is another constant $C_{T}$ such that, for all $\gamma\in{\mathcal{A}}$, and for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$:
@@
Without loss of generality, assume $y<y^{\prime}<y+1$. Then:
$L^{y^{\prime},\gamma}_{T}(\lambda)-L^{y,\gamma}_{T}(\lambda)\leq(e^{C_{T}(y^{\prime}-y)}-1)L^{y,\gamma}_{T}(\lambda)\leq C_{T}e^{C_{T}}|y-y^{\prime}|L^{y,\gamma}_{T}(\lambda).$
Similarly:
$L^{y,\gamma}_{T}(\lambda)-L^{y^{\prime},\gamma}_{T}(\lambda)\leq C_{T}e^{C_{T}}|y-y^{\prime}|L^{y^{\prime},\gamma}_{T}(\lambda).$
Combining these, the result follows. ∎
@@
Since $L^{y^{\prime},\gamma}_{T}(\lambda)\leq e^{C_{T}}L^{y,\gamma}_{T}(\lambda)$, there exists another constant $C_{T}$ such that, for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$,
@@
There exists a constant $C_{T}$ such that, for all $\gamma\in{\mathcal{A}}$, and for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$:
@@
The term $A$ satisfies:
$A\leq C_{T}|y-y^{\prime}|\mathbb{E}_{\mathbb{Q}}L^{\gamma,y}_{t,T}(\Lambda)\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})=C_{T}|y-y^{\prime}|J^{\gamma}_{2}(t,y,z,n).$
The term $B$ satisfies:
@@
Here, we used that $\sup_{n\in\mathbb{N}}\lVert\partial_{z}\Psi_{\mu}(n,\cdot)\rVert<\infty$. Combining these, the result follows. ∎
Combining these results, we deduce that:
@@
Let $y,y^{\prime}\in\mathbb{R}$ be fixed. Consider $\epsilon>0$ and an $\epsilon$-optimal control $\gamma$ such that:
@@
Letting $\epsilon\downarrow 0$, we obtain:
@@
Additionally, note that:
@@
By exchanging the roles of $y$ and $y^{\prime}$, we conclude:
@@ 3.1.2 Regularity with respect to $z$
The regularity with respect to $z$ is established using similar arguments. Recall that:
@@
There exist constants $C,\eta>0$ such that, for all $n\in\mathbb{N}$:
@@
    For all $z,z^{\prime}\geq 0$, $\lvert\Phi_{\mu}(n,z)-\Phi_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|\Phi_{\mu}(n,z)$.
@@
    For all $z,z^{\prime}\geq 0$, $\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|[m_{\mu}(n,z)+m_{\mu}(n,z^{\prime})]$.
@@
    For all $z,z^{\prime}\geq 0$ with $|z-z^{\prime}|<\eta$, $\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|m_{\mu}(n,z)$.
@@
For the first point:
@@
Since $Supp(\mu)\subset[0,\lambda_{\max}]$, the result follows. For the second point:
@@
Using the first point, we deduce that the second inequality holds. For the third point, we verify that, for $|z-z^{\prime}|\leq\eta$, with $\eta$ sufficiently small:
@@
Since $\lvert\Phi_{\mu}(n,z^{\prime})-\Phi_{\mu}(n,z)\rvert\leq\frac{1}{2}\Phi_{\mu}(n,z)$ for $\lvert z-z^{\prime}\rvert$ sufficiently small, the result follows. ∎
Since under $\mathbb{Q}$, the Poisson process $N$ is independent of $\Lambda$, we have:
@@
From the third point of the previous lemma, for all $z,z^{\prime}\geq 0$ with $\lvert z-z^{\prime}\rvert\leq\eta$, it follows that:
@@
Thus, as previously, we deduce:
@@
This completes the proof of Proposition 3.2.
@@ 3.2 A first dynamic programming principle
Recall that the value function is defined by:
$v(t,y,z,n)$ $:=\inf_{\gamma\in{\mathcal{A}}_{t}}\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\left[\int_{t}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+\kappa\Psi_{\mu}(n+N_{T}-N_{t},Z_{T}^{t,y,z,\gamma})\right]$
$=\inf_{\gamma\in{\mathcal{A}}_{t}}J^{\gamma}(t,y,z,n).$
@@
For all $s\in[t,T]$, it holds that:
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}_{t}}$ $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left[\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right].$
We now provide the proof of this result. Recall that for $\gamma\in{\mathcal{A}}_{t}$,
@@
We start with the following key lemma. Given $\gamma\in{\mathcal{A}}_{t}$, $\mathrm{w}\in\Omega$, and $s\geq t$, we set
$\gamma^{s,\mathrm{w}}:=\omega\mapsto\gamma(\mathrm{w}\otimes_{s}\omega),$
where for all $u\geq 0$:
$(\mathrm{w}\otimes_{s}\omega)(u)=\begin{cases}\mathrm{w}(u)\quad\text{ if }u\leq s\\ \mathrm{w}(s)+\omega(u)-\omega(s)\quad\text{ if }u>s.\end{cases}$
@@
Let $\gamma\in{\mathcal{A}}_{t}$ and let $S$ be a ${\mathcal{F}}^{N}$-stopping time such that $S\in[t,T]$. Then,$\mathbb{Q}(\mathrm{d}\omega)$ a.s.:
$\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\int_{S}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u\mid{\mathcal{F}}^{N}_{S}\right](\omega)$
$\quad=\langle L^{y,\gamma^{S(\omega),\omega}}_{t,S(\omega)},m_{\mu}(n,z)\rangle J^{\gamma^{S(\omega),\omega}}_{1}(S(\omega),Y^{t,y,\gamma^{S(\omega),\omega}}_{S(\omega)},Z^{t,y,z,\gamma^{S(\omega),\omega}}_{S(\omega)},n+N_{S(\omega)}-N_{t}),$
and:
$\mathbb{E}_{\mathbb{Q}}[\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})~{}|~{}{\mathcal{F}}^{N}_{S}](\omega)$
$\quad=\langle L^{y,\gamma^{S(\omega),\omega}}_{t,S(\omega)},m_{\mu}(n,z)\rangle J^{\gamma^{S(\omega),\omega}}_{2}(S(\omega),Y^{t,y,\gamma^{S(\omega),\omega}}_{S(\omega)},Z^{t,y,z,{\gamma^{S(\omega),\omega}}}_{S(\omega)},n+N_{S(\omega)}-N_{t}).$
@@
First, note that for all $\lambda\geq 0$:
$L^{y,\gamma}_{t,T}(\lambda)=L^{y,\gamma}_{t,S}(\lambda)L^{Y^{t,y,\gamma}_{S},\gamma}_{S,T}(\lambda).$
Combined with (2.9), this implies:
$\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\langle L^{Y^{t,y,\gamma}_{S},\gamma}_{S,T},m_{\mu}(n+N_{S}-N_{t},Z^{t,y,z,\gamma}_{S})\rangle=\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle.$
We now prove the second equality. Let $A\in{\mathcal{F}}^{N}_{S}$. To simplify notation, define:
$y^{\gamma}_{S}:=Y^{t,y,\gamma}_{S},\quad z^{\gamma}_{S}:=Z^{t,y,z,\gamma}_{S},\quad n_{S}:=n+N_{S}-N_{t}.$
Additionally, set:
$\Delta Z^{\gamma}_{S,T}:=Z_{T}^{S,y^{\gamma}_{S},z^{\gamma}_{S},\gamma}-z^{\gamma}_{S}=\int_{S}^{T}g(Y^{t,y,\gamma}_{u})\mathrm{d}u.$
Then:
@@
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\langle L^{y^{\gamma}_{S},\gamma}_{S,T},m_{\mu}(n_{S},z^{\gamma}_{S})\rangle\mathbbm{1}_{A}\Psi_{\mu}(n_{S}+N_{T}-N_{S},z^{\gamma}_{S}+\Delta Z^{\gamma}_{S,T})$
$\quad=\int\langle L^{y,\gamma}_{t,S}(\omega),m_{\mu}(n,z)\rangle\mathbbm{1}_{A}(\omega)$
$\qquad~{}~{}~{}~{}~{}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y^{\gamma}_{S},\gamma}_{S,T},m_{\mu}(n_{S},z^{\gamma}_{S})\rangle\Psi_{\mu}(n_{S}+N_{T}-N_{S},z^{\gamma}_{S}+\Delta Z^{\gamma}_{S,T})~{}|~{}{\mathcal{F}}^{N}_{S}\right](\omega)\mathbb{Q}(\mathrm{d}\omega)$
$\quad=\int\langle L^{y,\gamma}_{t,S}(\omega),m_{\mu}(n,z)\rangle\mathbbm{1}_{A}(\omega)J^{\gamma^{S(\omega),\omega}}_{2}(S,y^{\gamma^{S(\omega),\omega}}_{S},z^{\gamma^{S(\omega),\omega}}_{S},n_{S})(\omega)\mathbb{Q}(\mathrm{d}\omega).$
The last equality follows from the strong Markov property of the standard Poisson process. Since this holds for any $A\in{\mathcal{F}}^{N}_{S}$, the result is established. ∎
We now provide the proof of Proposition 3.9, which follows from standard arguments. Recall that:
$J^{\gamma}(t,y,z,n)=J^{\gamma}_{1}(t,y,z,n)+\kappa J^{\gamma}_{2}(t,y,z,n).$
Using Lemma 3.10 with a deterministic time $S=s\in[t,T]$, we have:
@@
It follows that:
$v(t,y,z,n)\geq\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left\{\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right\}\right].$
For the reverse inequality, fix $\epsilon>0$ and $s\in(t,T)$. Using the regularity of $J^{\gamma}$ and $v$ established in Section 3.1, there exists $\alpha>0$ such that, for all $n\in\mathbb{N}$, there is a partition $(B_{i}^{n})_{i\in\mathbb{N}}$ of $\mathbb{R}\times\mathbb{R}_{+}$ with centers $(y_{i}^{n},z_{i}^{n})\in B_{i}^{n}$ such that, for all $(n,i)\in\mathbb{N}^{2}$ and $(y,z)\in B_{i}^{n}$:
@@
For each $(y_{i}^{n},z_{i}^{n},n)$, there exists an $\epsilon$-optimal control $\alpha(y_{i}^{n},z_{i}^{n},n)\in{\mathcal{A}}$ such that:
@@
Define the control $\gamma\in{\mathcal{A}}$ by:
@@
where $\alpha^{i,n}:=\alpha(y_{i}^{n},z_{i}^{n},n)$ for $i,n\geq 0$ and $n_{s}:=n+N_{s}-N_{t}$. To simplify notation, let $y_{s}=Y^{t,y,\gamma}_{s}$ and $z_{s}=Z^{t,y,z,\gamma}_{s}$. Then:
@@
Letting $\epsilon\rightarrow 0$ and noting that $\gamma$ is arbitrary between $t$ and $s$, we deduce: